You decide to work part time at a local super market. The job pays $8.00 per hour and your work 21 hours a week. Your employed withholds 10 percent of your gross pay for federal taxes,5.65 percent for FICA taxes and 3 percent for state taxes.
What percentage of your gross pay is withheld for taxes?

Answer by stanbon(75887)

You decide to work part time at a local super market. The job pays $8.00 per hour and you work 21 hours a week. Your employer withholds 10 percent of your gross pay for federal taxes,5.65 percent for FICA taxes and 3 percent for state taxes.
What percentage of your gross pay is withheld for taxes?
-----
Ans: 10+5.65+3 = 10+8.65 = 18.65%
-------------
Cheers,
Stan H.
--------------